# Latent space models for grouped multiplex networks

### v1

@@ 2.2 Identifiability
Intuitively, Condition 2 ensures that, for each group $k$, there exist two individual components $U_{ks}$ and $U_{kt}$ whose columns are linearly independent of those in the “total shared” component $V,\ W_{k}]$. Condition 3 ensures that the shared component $V$ has linearly independent columns with those in the “total individual” components $W_{k_{1}},\ U_{k_{1}\ell_{1}}]$ and $W_{k_{2}},\ U_{k_{2}\ell_{2}}]$ for some groups $k_{1}$ and $k_{2}$.
3 Estimation
3.1 Likelihood maximization with nuclear norm penalization
Assuming the ground truth latent positions $V,\{W_{k}\}_{k=1}^{K},\{U_{k\ell}\}_{\mathcal{I}}$ satisfy the identifiability condition from Proposition 2.1, each of them is still identifiable only up to an indefinite orthogonal transformation. To avoid this issue, we propose to estimate their Gram matrices instead, and we denote them as follows:
$S=VI_{p_{0},q_{0}}V^{\top},$
$Q_{k}=W_{k}I_{p_{k},q_{k}}W_{k}^{\top},$ $k=1,\ldots,K,$ (7)
$R_{k\ell}=U_{k\ell}I_{p_{k\ell},q_{k\ell}}U_{k\ell}^{\top},$ $(k,\ell)\in\mathcal{I}.$
In practice, given the low-rank estimates $\hat{S},\hat{Q}_{k},$ and $\hat{R}_{k\ell}$, we can further extract $\hat{V},\hat{W}_{k},$ and $\hat{U}_{k\ell}$ with Adjacency Spectral Embedding (ASE) (RubinDelanchy2017). For example, to estimate $\hat{V}$, this method would suggest computing the truncated eigen-decomposition $\hat{S}=\hat{\bar{V}}\hat{\Gamma}\hat{\bar{V}}^{\top}$, where $\hat{\Gamma}\in\mathbb{R}^{\hat{d}_{0}\times\hat{d}_{0}}$ is a diagonal matrix containing the top $\hat{d}_{0}$ eigenvalues (in absolute magnitude), which are sorted on the diagonal in decreasing order with respect to their signed value, and $\hat{\bar{V}}\in\mathbb{R}^{n\times\hat{d}_{0}}$ contains the corresponding eigenvectors. In this case, $\hat{p}_{0}$ and $\hat{q}_{0}$ with $\hat{d}_{0}=\hat{p}_{0}+\hat{q}_{0}$ correspond to the number of positive and negative entries in $\hat{\Gamma}$, respectively. Then, we can set $\hat{V}=\hat{\bar{V}}|\hat{\Gamma}|^{1/2}$ where $|\cdot|$ takes absolute values of all entries in a matrix. Estimates $\hat{W}_{k}$ and $\hat{U}_{k\ell}$ can be computed similarly.
With the notations in (3.1) at hand, a natural way to fit the GroupMultiNeSS model would be to minimize the negative log-likelihood by exploiting the independence of adjacency matrix entries stated in (1):
$\mathcal{L}\left(S,\left\{Q_{k}\right\}_{k=1}^{K},\left\{R_{k\ell}\right\}_{\mathcal{I}};\ \left\{A_{k\ell}\right\}_{\mathcal{I}}\right)=\sum_{k=1}^{K}\mathcal{L}_{k}(S+Q_{k},\{R_{k\ell}\}_{\ell=1}^{m_{k}};\ \left\{A_{k\ell}\right\}_{\ell=1}^{m_{k}})$ (8)
where we denoted the negative likelihood of the $k$-th group by
$\mathcal{L}_{k}(S+Q_{k},\{R_{k\ell}\}_{\ell=1}^{m_{k}};\ \left\{A_{k\ell}\right\}_{\ell=1}^{m_{k}})=-\sum_{\ell=1}^{m_{k}}\sum_{i\leq j}\log f\left(A_{k\ell,ij};S_{ij}+Q_{k,ij}+R_{k\ell,ij},\phi\right)$ (9)
with the inner summation restricted to node pairs with $i<j$ in case the network is loop-free. Notice that the group log-likelihood $\mathcal{L}_{k}$ depends on variables $S+Q_{k}$ and $\{R_{k\ell}\}_{\ell=1}^{m_{k}}$ that appear only in the $k$-th group, giving a hint about a potential block structure in the optimization procedure. Indeed, we could first estimate $\widehat{S+Q_{k}}$ and $\{\hat{R}_{k\ell}\}_{\ell=1}^{m_{k}}$ within each group and then separate $\hat{S}$ from $\hat{Q}_{k},k=1,\ldots,K$ by optimizing the full likelihood $\mathcal{L}$ with all $\{\hat{R}_{k\ell}\}_{\mathcal{I}}$ fixed. An issue of direct optimization of (8) is that the parameters introduced in (3.1) are not restricted to be of low rank and can violate the required rank constraints:
$\begin{array}[]{llll}&\operatorname{rank}^{+}(S)=p_{0},&\operatorname{rank}^{-}(S)=q_{0}&\\ &\operatorname{rank}^{+}(Q_{k})=p_{k},&\operatorname{rank}^{-}(Q_{k})=q_{k}&\quad k=1,\ldots,K,\\ &\operatorname{rank}^{+}(R_{k\ell})=p_{k\ell},&\operatorname{rank}^{-}(R_{k\ell})=q_{k\ell}&\quad(k,\ell)\in\mathcal{I},\end{array}$
where $\operatorname{rank}^{+}(Z)$ and $\operatorname{rank}^{-}(Z)$ denote respectively the number of strictly positive and strictly negative eigenvalues of a symmetric matrix $Z$. Unfortunately, even in an oracle scenario where the dimensions of latent components are known, including the rank constraints in the optimization makes the problem non-convex and usually intractable. Instead, we can perform convex relaxation using a nuclear penalty regularization, resulting in the following two-stage optimization problem, where in the first stage, for each $k=1,\ldots,K$, we estimate $\widehat{S+Q_{k}},\{\hat{R}_{k\ell}\}_{\ell=1}^{m_{k}}$ by solving the following optimization problem:
@@ 3.1 Likelihood maximization with nuclear norm penalization
and in the second stage, we get $\hat{S},\{\hat{Q}_{k}\}_{k=1}^{K}$ by solving:
@@
Here, $\|\cdot\|_{*}$ denotes the nuclear norm of a matrix equal to the sum of its singular values, and $\{\lambda_{1k}\}_{k=1}^{K},\lambda_{2},\{\alpha_{2k}\}_{k=1}^{K},\{\alpha_{1k\ell}\}_{\mathcal{I}}$ are tunable hyperparameters (see details on tuning in Section 3.2). Since the nuclear norm is convex, these optimization problems are convex if the edge distribution density $f(\cdot;\theta,\phi)$ is log-concave in $\theta$.
@@
The formulated two-stage fitting procedure can be thought of as a bottom-to-top pruning of the GroupMultiNeSS tree in Figure 1. Indeed, estimation of the individual components in the first stage can be interpreted as “cutting off” the leaves from the tree, since they are no longer updated in the second stage. Similarly, the second-stage separation of the group and shared components can be interpreted as cutting off the tree’s second-level nodes, which act as new leaves after the first-stage completion. This analogy suggests a natural extension of the proposed fitting procedure to the model, which generalizes GroupMultiNeSS by allowing the tree structure of the network layers to involve a more complex hierarchy, for example, by including additional subgroup latent components.
To minimize the loss functions in (10) and (11), we perform a coordinate descent by iteratively updating one of the optimized matrices while keeping the others fixed. In turn, each matrix update can be performed using a proximal gradient method, specifically to optimize a loss function $L$ combined with the non-differentiable nuclear norm penalty (FithianMazumder2018). Given the gradient step-size $\eta>0$, the update at iteration $t\geq 1$ is performed as follows
$Z^{(t+1)}=\operatorname*{argmin}_{Z\in\mathbb{R}^{n\times n}}\Bigl\{{1\over 2\eta}\Bigl\|Z-\Bigl[Z^{(t)}-\eta\nabla L(Z^{(t)})\Bigr]\Bigr\|_{F}^{2}+\rho\|Z\|_{*}\Bigr\}$ (12)
where $L$, $Z$, and $\rho$ are place-holders respectively, for a loss function, an optimized parameter matrix, and a hyperparameter associated with the nuclear norm penalty. In the first stage, the loss is $\mathcal{L}_{k}$ and the matrices are $S+Q_{k}$ and $R_{k\ell}$ with respective thresholds $\lambda_{1k}$ and $\lambda_{1k}\alpha_{1k\ell}$. In the second stage, the loss is $\mathcal{L}$ and the matrices are $S$ and $Q_{k}$ with thresholds $\lambda_{2}$ and $\lambda_{2}\alpha_{2k}$, respectively. It is well-known that the solution to the optimization problem in (12) has an explicit form in terms of the so-called soft-thresholding operator
$Z^{(t+1)}=\mathcal{T}_{\eta\rho}\left(Z^{(t)}-\eta\nabla L(Z^{(t)})\right).$ (13)
For a diagonal matrix $\Sigma\in\mathbb{R}^{n\times n}$, the soft-thresholding operator $\mathcal{T}_{s}$ with parameter $s$ is defined as
$\mathcal{T}_{s}(\Sigma)=\operatorname{diag}[(\Sigma_{11}-s)_{+},\ldots,(\Sigma_{nn}-s)_{+}]$
with $(x)_{+}:=\max\{0,x\}$ and for a non-diagonal matrix with singular value decomposition $Z=U\Sigma V^{\top}$ as
$\mathcal{T}_{s}(Z)=U\mathcal{T}_{s}(\Sigma)V^{\top}.$
Intuitively, the soft-thresholding operator in the update rule (13) mimics a projector onto the space of low-rank matrices by truncating the singular values of a parameter matrix after each gradient step. In case the ground-truth rank $d$ of $Z$ was known, soft-thresholding could be replaced by hard-thresholding
@@
where
$Z]_{d}:=\operatorname*{argmin}_{\operatorname{rank}\left(Z^{\prime}\right)\leq d}\left\|Z-Z^{\prime}\right\|_{F}$ (15)
is well-defined by the Eckart-Young Theorem as the truncation of the SVD of $Z$ to the largest $d$ singular values.
With the derived update rule for each parameter matrix at hand, we can define the end-to-end optimization procedure as Algorithm 1. In the first stage of the algorithm, we solve Problem (10) for every $k=1,\ldots,K$ by alternating the updates of $S+Q_{k}$ and $\{R_{k\ell}\}_{l=1}^{m_{k}}$ until convergence. In the second stage, we solve Problem (11) by alternating the updates of $S$ and $\{Q_{k}\}_{k=1}^{K}$. Each matrix update is performed according to (13) but with stage-specific learning rate $\eta_{1},\eta_{2}>0$ and parameter-specific normalization for convergence stability. In the first stage, we normalize $\eta_{1}$ by $m_{k}$ for updates of $S+Q_{k}$ and in the second stage, we normalize $\eta_{2}$ by $M$ and $m_{k}$ for updates of $S$ and $Q_{k}$, respectively, that is, we divide by the number of times these parameters appear in the joint log-likelihood (8). The parameter matrices obtained by running the alternating updates within each optimization subproblem are further passed to the optional refitting step aimed to debias the learned non-zero eigenvalues of the fitted matrices and so compensate for their excessive truncation caused by the nuclear norm penalization (see details in Section 3.4). The first and second stage refitting procedures will exhibit some principal differences, arising from the fact that the individual components $\hat{R}_{k\ell}$ fitted in the first stage should be kept fixed during the second stage refitting. Therefore, we denote the corresponding algorithms as $\it{FirstStageRefit}$ and $\it{SecondStageRefit}$, respectively.
For now, we do not explicitly state how to initialize parameters in both stages and denote the corresponding procedures for the first and second stages as $\it{FirstStageInit}$ and $\it{SecondStageInit}$, respectively. The detailed investigation of initialization options is postponed to Section 3.3. Notice that for initializing $S+Q_{k}$ in the first stage, we only use the $k$-th group layers $\{A_{k\ell}\}_{\ell=1}^{m_{k}}$ as $S+Q_{k}$ is essentially an independent variable that does not appear in any other group except for the group $k$ itself and so other layers cannot give extra information on its possible value.
@@
The computational bottleneck within every iteration of the first and second stages of Algorithm 1 is the computation of the SVD for each updated parameter matrix, which is needed for soft thresholding. If we computed a full SVD, each iteration within group $k$ of the first stage would require $O(m_{k}n^{3})$ operations, and each iteration within the second stage would take $O(Kn^{3})$. In addition to the discussed parallelization of the first-stage subproblems, we speed up the computations by using a truncated SVD for matrix update, which only finds its leading $r\ll n$ singular values, reducing the complexity to $O(m_{k}n^{2}r)$ and $O(Kn^{2}r)$, respectively. In our implementation, we use $r=\lceil\sqrt{n}\ \rceil$ as the default value.
@@
Algorithm 1 allows for many similar optimization alternatives, such as updating all three types of components ($S$, $Q_{k}$, $R_{k\ell}$) in a single loop, substituting the second stage loss from the negative log-likelihood to the quadratic loss between $S+Q_{k}$ and $\widehat{S+Q_{k}}$, or pre-estimating component ranks $\{d_{k}\}_{k=0}^{K}$ and $\{d_{k\ell}\}_{\mathcal{I}}$ first and then substituting the soft-thresholding updates by the hard-thresholding ones as described in (14). The discussion and comparison of these and other alternatives is postponed to Section A of the Appendix.
Notice that the optimization procedure suggested for the first-stage problem (10) essentially applies the fitting procedure of the MultiNeSS model to the layers $\{A_{k\ell}\}_{\ell=1}^{m_{k}}$ (as defined in Equation 5 of multiness) aiming to separate the individual components $\{R_{k\ell}\}_{\ell=1}^{m_{k}}$ from the “total shared” component $S+Q_{k}$. In contrast, the second-stage optimization is no longer so interpretable since we fixed the estimates of the individual components. To build some intuition about what happens in the second stage of Algorithm 1, we explicitly state its relationship to the MultiNeSS fitting procedure in the special case of the Gaussian edge-entry distribution. Intuitively, we demonstrate that in the Gaussian case, the second stage of the GroupMultiNeSS essentially fits the MultiNeSS model to the group-wise averaged residuals between the layers and their estimated individual components. The proof of this proposition can be found in Section B.2 of the Appendix.

### v2

@@ 2.2 Identifiability
Intuitively, Condition 2 requires that in each group $k$ there are at least two individual components with columns linearly independent of those shared by the whole group ($V,\ W_{k}]$) – if that is not the case, then the individual components still contain some common group information. Similarly, Condition 3 requires that the shared component $V$ has linearly independent columns from at least some $W_{k_{1}},\ U_{k_{1}\ell_{1}}]$ and $W_{k_{2}},\ U_{k_{2}\ell_{2}}]$, as otherwise not all shared information would have been separated.
3 Fitting the GroupMultiNeSS model
The true latent positions $V,\{W_{k}\}_{k=1}^{K},\{U_{k\ell}\}_{\mathcal{I}}$ are only identifiable up to a rotation. To avoid making an arbitrary choice of the rotation, we instead focus on estimating the Gram matrices
$S=VI_{p_{0},q_{0}}V^{\top},\ \ Q_{k}=W_{k}I_{p_{k},q_{k}}W_{k}^{\top},\ \ R_{k\ell}=U_{k\ell}I_{p_{k\ell},q_{k\ell}}U_{k\ell}^{\top},$ (7)
for $k=1,\ldots,K$, $(k,\ell)\in\mathcal{I}$. If needed, we can always extract estimates of the latent positions themselves by performing an SVD on the estimated Gram matrices, and fix a rotation to align latent positions across groups.
3.1 Likelihood maximization with a nuclear norm penalty
A natural way to estimate parameters (7) is to minimize the negative log-likelihood under the GroupMultiNeSS model, which is easy to write down due to independence. We can write the full data likelihood as
$\mathcal{L}\left(S,\left\{Q_{k}\right\}_{k=1}^{K},\left\{R_{k\ell}\right\}_{\mathcal{I}};\ \left\{A_{k\ell}\right\}_{\mathcal{I}}\right)=\sum_{k=1}^{K}\mathcal{L}_{k}(S+Q_{k},\{R_{k\ell}\}_{\ell=1}^{m_{k}};\ \left\{A_{k\ell}\right\}_{\ell=1}^{m_{k}})\,$, (8)
where
$\mathcal{L}_{k}(S+Q_{k},\{R_{k\ell}\}_{\ell=1}^{m_{k}};\ \left\{A_{k\ell}\right\}_{\ell=1}^{m_{k}})=-\sum_{\ell=1}^{m_{k}}\sum_{i\leq j}\log f\left(A_{k\ell,ij};S_{ij}+Q_{k,ij}+R_{k\ell,ij},\phi\right).$ (9)
The sum over $i\leq j$ can be replaced with the sum over $i<j$ if we do not want to model diagonal entries.
Imposing low-rank constraints on the matrices defined (7) turns the optimization problem (8) non-convex and typically intractable. Instead, we follow the common strategy of replacing the rank constraints with a nuclear norm penalty, which gives us a convex optimization problem as long as the edge distribution density $f(\cdot;\theta,\phi)$ is log-concave in $\theta$.
Note that the group log-likelihood $\mathcal{L}_{k}$ depends on $S+Q_{k}$ and $\{R_{k\ell}\}_{\ell=1}^{m_{k}}$ for that $k$ only, allowing for a two-stage approach to solving the optimization problem: we can first estimate $\widehat{S+Q_{k}}$ and $\{\hat{R}_{k\ell}\}_{\ell=1}^{m_{k}}$ within each group and then estimate $\hat{S}$ and $\hat{Q}_{k},k=1,\ldots,K$ by optimizing the full likelihood $\mathcal{L}$ with all $\{\hat{R}_{k\ell}\}_{\mathcal{I}}$ fixed. Formally, in the first stage we solve the optimization problem
@@ 3.1 Likelihood maximization with a nuclear norm penalty
and in the second stage, we solve
@@
Here, $\|\cdot\|_{*}$ denotes the matrix nuclear norm (the sum of its singular values), and $\{\lambda_{1k}\}_{k=1}^{K},\lambda_{2}$, $\{\alpha_{2k}\}_{k=1}^{K},\{\alpha_{1k\ell}\}_{\mathcal{I}}$ are hyperparameters; choosing them in practice is discussed in Section 3.2).
@@
The two-stage fitting procedure can be thought of as a bottom-to-top fitting of the GroupMultiNeSS tree in Figure 1: the first stage estimates the "leaves" at the bottom, which stay fixed after, and the second stage estimates the next level of the tree. This procedure is thus easily extended to nested groups, as long as they are arranged in a hierarchical tree.
We solve the optimization problems (10) and (11) by block coordinate descent, updating one of the matrices at a time while keeping the others fixed. Each matrix update is performed via a proximal gradient method developed for the nuclear norm penalty by (Fithian and Mazumder, 2018). Given the gradient step size $\eta>0$, the update at iteration $t\geq 1$ is given by
$Z^{(t+1)}=\operatorname*{argmin}_{Z\in\mathbb{R}^{n\times n}}\Bigl\{{1\over 2\eta}\Bigl\|Z-\Bigl[Z^{(t)}-\eta\nabla L(Z^{(t)})\Bigr]\Bigr\|_{F}^{2}+\rho\|Z\|_{*}\Bigr\}\ ,$ (12)
where $L$, $Z$, and $\rho$ stand for generic loss function, matrix being optimized over, and the nuclear norm penalty tuning parameter, respectively. It is well-known that the solution to the optimization problem in (12) is given by the soft-thresholding operator,
$Z^{(t+1)}=\mathcal{T}_{\eta\rho}\left(Z^{(t)}-\eta\nabla L(Z^{(t)})\right),$ (13)
where for a matrix with the singular value decomposition $Z=U\Sigma V^{\top}$ the soft-thresholding operator $\mathcal{T}_{s}$ is defined as
$\mathcal{T}_{s}(Z)=U\operatorname{diag}[(\Sigma_{11}-s)_{+},\ldots,(\Sigma_{nn}-s)_{+}]V^{\top}\,$,
with $(x)_{+}:=\max\{0,x\}$. Intuitively, the soft-thresholding operator projects the previous iterate onto the space of low-rank matrices by truncating its singular values.
If the true rank $d$ of $Z$ is known (oracle estimator), soft thresholding is usually replaced by hard thresholding,
@@
where $Z]_{d}:=\operatorname*{argmin}_{\operatorname{rank}\left(Z^{\prime}\right)\leq d}\left\|Z-Z^{\prime}\right\|_{F}$, by the Eckart-Young theorem, is the truncation of the SVD of $Z$ to the largest $d$ singular values.
We summarize the entire optimization procedure in Algorithm 1. In the first stage, we solve problem (10) for every $k=1,\ldots,K$ by alternating between updates of $S+Q_{k}$ and $\{R_{k\ell}\}_{l=1}^{m_{k}}$ until convergence. In the second stage, we solve problem (11) by alternating between updates of $S$ and $\{Q_{k}\}_{k=1}^{K}$. Each matrix update is performed according to (13) but with stage-specific learning rates $\eta_{1},\eta_{2}>0$ and parameter-specific normalization for stability: we normalize $\eta_{1}$ by $m_{k}$ for updates of $S+Q_{k}$ and $\eta_{2}$ by $M$ and $m_{k}$ for updates of $S$ and $Q_{k}$, respectively, that is, we divide by the number of times these parameters appear in the joint log-likelihood (8). The parameter matrices obtained by running the alternating updates within each optimization subproblem are passed to the optional refitting step, a common post-processing step for the spectral regularization methods (Mazumder et al., 2010). This procedure “unshrinks” non-zero eigenvalues of the fitted matrices to compensate for excessive truncation caused by the nuclear norm penalty (see details in Section A of the Appendix). The first and second stage refitting procedures differ because the individual components $\hat{R}_{k\ell}$ from the first stage are kept fixed in the second stage; we refer to the algorithms as FirstStageRefit and SecondStageRefit, respectively.
We postpone the discussion of initialization to Section 3.3 and for now state the algorithm for generic initialization algorithms FirstStageInit and SecondStageInit.
@@
The main computational bottleneck within every iteration of the first and second stages of Algorithm 1 is computing the SVD for each updated parameter matrix, which is needed for soft thresholding. Computing the truncated SVD for $r$ leading singular values gives the total iteration complexity of $O(m_{k}rn^{2})$ for Stage I and $O(Krn^{2})$ for Stage II. In our implementation, we take $r=\sqrt{n}$ as the default value. We also parallelize the first stage computations across groups.
@@
One alternative approach to optimizing the joint log-likelihood in (8) would be to add nuclear norm penalties for all of the matrices $S,Q_{k},R_{k\ell}$ and update all of them in a single loop. An important advantage of Algorithm 1 over this is that the loop over groups $k=1,\ldots,K$ can be parallelized. While we then have to perform a second stage optimization, we expect it to be much less computationally expensive than each first-stage problem, because the number of parameter matrices in the second stage is $K+1$, which is typically much smaller than $m_{k}+1,k=1,\ldots,K$ for each of the parallelized optimizations in the first stage. Empirically, we indeed observed that this approach is not only significantly more expensive in terms of CPU time and the number of SVD computations, but also less stable in terms of convergence.
Remark 5.
Another alternative is to apply a non-convex approach analogous to that of Tian et al. (2024), which pre-estimates the latent component ranks via the Shared Space Hunting (SSH) approach; we could apply SSH and then replace soft with hard thresholding. The key advantage of this approach is avoiding cross-validation, and estimating ranks independently of the assumed likelihood. However, SSH is only guaranteed to accurately estimate ranks that are much smaller than $n$, and we have empirically observed it is often unstable in harder settings. When the ranks are misspecified, our Algorithm 1 performs significantly better, and it gives comparable or marginally worse results when ranks are estimated correctly. While both are valid approaches, we have opted to use cross-validation for stability, in spite of its higher computational cost.
The optimization procedure in the first stage (10) essentially applies the MultiNeSS fitting algorithm to layers $\{A_{k\ell}\}_{\ell=1}^{m_{k}}$ (as defined in Equation 5 of MacDonald et al. (2022)), aiming to separate the individual components $\{R_{k\ell}\}_{\ell=1}^{m_{k}}$ from the $S+Q_{k}$, the shared component for group $k$. In contrast, the second-stage optimization is not directly interpretable as an application of MultiNeSS, since we fix the estimated individual components. To provide some intuition about the second stage of Algorithm 1, we explicitly state its relationship to the MultiNeSS algorithm in the special case of the Gaussian edge-entry distribution, showing that it essentially fits the MultiNeSS model to the group-wise averaged residuals between the layers and their estimated individual components. The proof of this proposition can be found in Section B.2 of the Appendix.